OpenStudy (anonymous):

Identify the discontinuities, and find the zeros of each rational function. f(x) = x^2-100/x+10

OpenStudy (anonymous):

\[f(x)={x^2-100}{x+10}=\frac{(x-10)(x+10)}{x+10}=x-10,x\neq -10\] the discontinuity is at \(x=-10\) and the zero is at \(x=10\)

OpenStudy (anonymous):

thanks